OpenStudy (anonymous):

Cara is making punch for a party. She wants to combine 20% real fruit juice drink with a 100% real fruit juice drink to get 16 cups of a mixture that is 80% real fruit juice. How much of each type of drink should she use?

OpenStudy (anonymous):

lets put \(x\) as the amount of 100% juice then the amount of 20% juice is \(16-x\) since the total is 16

OpenStudy (anonymous):

you want to have 80% of 16 = 12.8 real juice, so solve \[x+.2(16-x)=12.8\] for \(x\)

OpenStudy (anonymous):

probably easier to multiply by 10 and solve \[10x+2(16-x)=128\]

OpenStudy (anonymous):

\[10x+32-2x=128\] \[8x+32=128\] two more steps and you are done

OpenStudy (anonymous):

x would equal 12

OpenStudy (anonymous):

yes, so 12 cups of real juice and the rest (4 cups) or 20% juice

OpenStudy (anonymous):

Thank you satellite73!